I need the answer to include solver and excel. Thank you.A Company produces two products. Relevant information for each product is shown in the Table below. The company has a goal of $48 in profits and incurs $1 penalty for each dollar it falls short of this goal. A total of 32 hours of labor are available. A $2 penalty is incurred for each hour of overtime (labor over 32 hours) used, and $1 penalty is incurred for each hour of available labor that is unused. Marketing considerations require at least 5 units of product 1 and 10 units of product 2 be produced. For each unit of either product by which production falls short of demand, a penalty of $5 is assessed Labor Required Contribution to profit Product 1 4 hours $4 Product 2 2 hours $2 Formulate a weighted goal programming that can be used to minimize the penalty incurred by the company. Do NOT solve, just formulate. 1. 2. Suppose that company sets (in order of importance) the following goals: a. Goal 1: Avoid underutilization of labor b. Goal 2: Meet demand for product 1 c. Goal 3: Meet demand for product 2 d. Goal 4: Do not use overtime Formulate and solve the preemptive goal programming model for this situation using Excel solver. Describe clearly the optimal solution to this problem using a managerial statement.
Answer #1

x1: Number of units of product 1 to produce.

            x2: Number of units of product 2 to produce.

Product 1, x1

Product 2, x2

Available

Penalty

Labor:

4 Hours

2 Hours

32 Hours

$2 / overtime

$1 / unused

Profit:

$4

$2

$48

$1 / dollar short

Requirement:

At least 7 units

At least 10 units

Penalty:

$5 / shortage

$5 / shortage

Constraints:

            4x1 + 2x2 + S1- - S1+ = 32

            x1 + S2- - S2+ = 7

            x2 + S3- - S3+ = 10

Quant Input Input Data of The Problem 0784N09a MIN +0 MIN +0 MIN +0 MIN +0 Subject to Page: .1 SIP +0 X1 +0 S2P +0 X1 +0 S2P +0 X1 +0 X2+0 X2+0 X2+0 S3P SIN+0 S3P SIN +0 S3P S2N S1P +1.00000S2N S2N S2N S1P +0 S2P +0 S3P (1) +4.00000X1 +2.00000X2 +1.00000SIN -1.00000S1P+0 (2) 1.00000Xi +0 (3) +0 S2N S2P + S3N +0 X2+0 S3N +0 S3P +32.0000 S1N +0 S3P = +7.00000 SIN +0 S1P +1.00000S2N S1P +0 S2N S2P +1. 00000S3N-1.00000S3P +10.0000 Quant Output: Summarized Solution for 0784NO9a Page 1 OpportunitylOpportunity l OpportunitylOpportunityl Number I Variable Solution ICost-obj. 11Cost-obj. 21Cost-obj 31 Cost-obj. 41 I+7.0000000 1 +10.000000 2 I SIN I S1P I S2N I S2P 0 1.0000000 이 +1 . 00000001 1 +16.000000 이 +1 . 00000001 01-4 . 00000001 01 +4.0000000 7 0 l +1 . 00000001-2 . 00000001 01 +2.00000001 Priority Level 1: Minimized Objective Function (Goal) = Priority Level 2: Minimized Objective Function (Goal) = Priority Level 3: Minimized Objective Function (Goal) = Priority Level 4: Minimized Objective Function (Goal) = +16.000000 iteration = 11 Elapsed CPU second = 0
Answer #2
Im need the answer
answered by: Othman Warde